[figure: 9.7 cm]

THE MAGNETIC BEHAVIOUR OF (Y_{1−x}Tb_{x})_{3}Co_{11}B_{4} INTERMETALLIC COMPOUNDS

E. Burzo, V. Pop, I.Gr. Deac, R. Tetean Babes-Bolyai University, Faculty of Physics, RO 3400 Cluj-Napoca, Romania

The R_{3}Co_{11}B_{4} compounds belong to R_{n+1}Co_{3n+5}B_{2n} intermetallic systems. The R_{3}Co_{11}B_{4} borides crystallize in a hexagonal structure of Ce_{3}Co_{11}B_{4} type having the space group P 6/mmm. The composition dependencies of the lattice parameters in (Y_{1−x}T b_{x})_{3}Co_{11}B_{4} compounds is linearly and opposite for a and c parameters. The magnetic properties of these compounds were investigated in the temperature range 5 - 1000 K and external fields up to 9 T. The Y_{3}Co_{11}B_{4} compound is ferromagnetically ordered. For the compounds with x > 0 a ferrimagnetic ordering is evidenced. The Curie temperatures increase when yttrium is substituted by terbium, as a consequence of the increasing of the exchange interactions in the system. The magnetic behaviour of cobalt in the above system is analyzed.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : R.Tetean

Address for correspondence : Babes-Bolyai University, Faculty of Physics, RO-3400 Cluj-Napoca, Romania Tel: ++4064194315 ext: 196, Fax: ++4064191906

Email address : rote@phys.ubbcluj.ro